OpenStudy (anonymous):

factor y=x^4-8x^2+16

OpenStudy (callisto):

y=x^4-8x^2+16 = (x^2 - 4) ^2 = (x-2)^2 (x+2)^2

OpenStudy (anonymous):

urs is actally confusing to be honest, wish i can take bck the medal.